A box moves 10 m horizontally when force F=20 N is applied at an angle θ=120°. What is


the work done on the box by the force during the displacement?

1
Expert's answer
2022-05-28T08:22:16-0400
"W=Fd\cos\theta\\
=\rm 20\: N\times 10\: m\times \cos120^\circ=-100\:J"